Fill in the blank. Suppose that the average starting salary for student with an Economics degree is $61,968.88 with a standard deviation of $5,909.25. A random sample of 163 recent graduates with a job is taken. There is a 42% chance that the average salary is greater than $ ________.

1) 60,775.84

2) There is not enough information to determine this.

3) 61,875.43

4) 62,062.33

5) 63,161.92
Answer #1

Given that,

mean = \mu = 61968.88

standard deviation = \sigma = 5909.25

\sigma\bar x=\sigma /\sqrtn=5909.25/\sqrt163=462.848

Using standard normal table,

1-P(Z < z) =42 %

= P(Z < z) = 1-0.42  

= P(Z < 0.20) = 0.58

z = 0.20

Using z-score formula  

x= z * \sigma + \mu

Using z-score formula  

\bar x = z * \sigma \bar x+ \mu \bar x

x= 0.20*462.848+61968.88

x= 62062.33 (rounded close)